OpenStudy (anonymous):

3x√8x -4√x^3 +3x√72x x=2

OpenStudy (agent0smith):

\[\Large 3x \sqrt {8x} -4 \sqrt {x^3} +3x \sqrt {72x}\]when x=2? If that's it, replace all the x's with 2, and simplify.

OpenStudy (anonymous):

Could you work it out for me, please? I am really bad at math.

OpenStudy (agent0smith):

I'll plug in the 2's for you, i'm sure you can simplify it somewhat from here\[\Large 3*2 \sqrt {8*2} -4 \sqrt {2^3} +3*2 \sqrt {72*2}\]

OpenStudy (anonymous):

6√16 -4√8 +6√144 right?

OpenStudy (agent0smith):

Correct. Now two of those square roots are perfect squares which are easy to simplify...

OpenStudy (anonymous):

6√4 -4√8 +6√12 ?

OpenStudy (agent0smith):

The square root of 16 is equal to 4, not the square root of 4. The square root of 144 is equal to 12, not the square root of 12. \[\large \sqrt {16} = 4\]\[\large \sqrt {144} =12\]

OpenStudy (anonymous):

How would you write that with the equation?

OpenStudy (agent0smith):

Replace the sqrt of 16 with 4, and the sqrt of 144 with 12...remember that the numbers in front are multiplying like so \[\Large 6*\sqrt{16} -4 \sqrt 8 +6 *\sqrt {144 } = \]

OpenStudy (anonymous):

So it would be 6*4 -4√8 +6*12 ?

OpenStudy (agent0smith):

Yes. So simplify that as much as you can, then we'll deal with the last sqrt.

OpenStudy (agent0smith):

Remember you can change the order... 6*4 -4√8 +6*12 = 6*4 + 6*12 - 4√8

OpenStudy (anonymous):

24 -4√8 +72

OpenStudy (anonymous):

96 -4√8

OpenStudy (agent0smith):

Now simplify the sqrt 8 like so.... since 8 is 4*2 \[\Large 96 -4 \sqrt 8 = 96 - 4 \sqrt { 4*2}\] \[\Large 96 - 4* \sqrt 4 \sqrt 2\]

OpenStudy (anonymous):

Would it make it 218.81?

OpenStudy (agent0smith):

No. You're keeping exact values. Simplify what i just posted.

OpenStudy (anonymous):

Not sure if I know how to do that. Sorry.

OpenStudy (agent0smith):

You just did it with the other square roots above, now try it with this. Simplify everything that you can, leave any square roots that you can't. \[\Large 96 - 4* \sqrt 4 * \sqrt 2\]

OpenStudy (anonymous):

96-4 * 2√2*√2

OpenStudy (agent0smith):

The square root of 4 is 2, so replace the sqrt of 4 with 2...\[\Large 96 - 4* \sqrt 4 * \sqrt 2 = \] \[\Large 96 - 4* 2* \sqrt 2\]Now you can finish simplifying. There's only the 4 and 2 to multiply together.

OpenStudy (agent0smith):

And you're done :)

OpenStudy (anonymous):

Thank you so much!

OpenStudy (agent0smith):

Welcome :)